A company used straight-line depreciation for an item of equipment that cost $18,250, had a salvage value of $4,600 and a six-year useful life. After depreciating the asset for three complete years, the salvage value was reduced to $1,825 but its total useful life remained the same. Determine the amount of depreciation to be charged against the equipment during each of the remaining years of its useful life Multiple Choice $1,365. $4140 $4,600 $7008. $3,200.
Answer #1

Correct answer--------------$3,200

Working

Straight line Method
A Cost $ 18,250
B Residual Value $ 4,600
C=A - B Depreciable base $ 13,650
D Life [in years left ]                                  6
E=C/D Annual SLM depreciation $ 2,275.00

.

Depreciation schedule-Straight line method
Year Book Value Depreciation expense Accumulated Depreciation Ending Book Value
1 $          18,250.00 $               2,275.00 $     2,275.00 $     15,975.00
2 $          15,975.00 $               2,275.00 $     4,550.00 $     13,700.00
3 $          13,700.00 $               2,275.00 $     6,825.00 $     11,425.00

.

Straight line Method
A Book value of year 3 end $ 11,425
B Residual Value $ 1,825
C=A - B Depreciable base $ 9,600
D Life [in years left ]                                  3
E=C/D Annual SLM depreciation $ 3,200.00